Project

General

Profile

NDOS Commissioning Meeting 2011/08/08

Attendance: Jon, Chad, Mark Mathis, Kurt, Martin, Gary, Mark Messier, Rafael, Leon, Athans,

Agenda:

  1. Beam & Detector Update
  2. DAQ Status & Plans
  3. DCS Status & Plans
  4. AOB

Beam & NDOS Update

  • Over the weekend, Denis did some TECC-noise tests on the detector. Did DSO scans prior to cooling entire detector and then again after the tests. Afterward, 5 APDs were "lost". DSO scans done after seem to have recovered these. Denis is analyzing the data collected.
  • Last time we were able to make a Booster timing peak was April. Most recent peak we tried to make was the end of July. Chad will try to make peak for last few weeks of running.
  • Leon:
    • still don't have ability to run overnight tests
    • other tests have been going on; 27 out of ~100 APDs have passed these tests. These are ones that were removed due to installation errors, dirt, etc., but are not ones that were found to have water on them.

DAQ Status & Plans

  • Peter has a script which he runs every few days that compares MINOS time of spill to hours. Last time he ran it looked fine. He will try to make this something that a shifter can use to check our timing.
  • Leon: clock and trigger are coming from the same source.
  • Peter/Kurt: new DCM firmware has some incompatibilities which are being addressed. Once the firmware is ready to go and tested, we may wish to deploy on the detector.

DCS Status & Plans

  • Nothing new to report.
  • Martin: would like to deploy new version of IOC.

AOB

  • None